# SR-CACO-2: A Dataset for Confocal Fluorescence Microscopy Image Super-Resolution

### v1

@@ Abstract
Confocal fluorescence microscopy is one of the most accessible and widely used imaging techniques for the study of biological processes at the cellular and subcellular levels. Scanning confocal microscopy allows the capture of high-quality images from thick three-dimensional (3D) samples, yet suffers from well-known limitations such as photobleaching and phototoxicity of specimens caused by intense light exposure, which limits its use in some applications, especially for living cells. Cellular damage can be alleviated by changing imaging parameters to reduce light exposure, often at the expense of image quality. Machine/deep learning methods for single-image super-resolution (SISR) can be applied to restore image quality by upscaling lower-resolution (LR) images to produce high-resolution images (HR). These SISR methods have been successfully applied to photo-realistic images due partly to the abundance of publicly available data. In contrast, the lack of publicly available data partly limits their application and success in scanning confocal microscopy. In this paper, we introduce a large scanning confocal microscopy dataset named SR-CACO-2 that is comprised of low- and high-resolution image pairs marked for three different fluorescent markers. It allows the evaluation of performance of SISR methods on three different upscaling levels (X2, X4, X8). SR-CACO-2 contains the human epithelial cell line Caco-2 (ATCC HTB-37), and it is composed of 22 tiles that have been translated in the form of 9,937 image patches for experiments with SISR methods. Given the new SR-CACO-2 dataset, we also provide benchmarking results for 15 state-of-the-art methods that are representative of the main SISR families. Results show that these methods have limited success in producing high-resolution textures, indicating that SR-CACO-2 represents a challenging problem. Our dataset, code and pretrained weights are available: https://github.com/sbelharbi/sr-caco-2.
@@ 1 Introduction
Deep learning methods have recently provided significant improvements in diverse image analysis tasks [18] and biosciences [31, 50]. In particular, it has allowed for considerable advances in single-image super-resolution (SISR) [70] where high-resolution (HR) images are restored from low-resolution (LR) photo-realistic images [38, 65, 73, 74]. However, the success of these methods relies heavily on the availability of large-scale public datasets to train a deep SISR model. Few works in microscopy imaging aim to leverage SISR [8, 11, 26, 37, 43, 46, 47]. This extends to different microscopy techniques [60] including standard fluorescence microscopy such as confocal [57], Structured Illumination Microscopy (SIM) [19], STimulated Emission Depletion (STED) microscopy [22], Single-Molecule Localization Microscopy (SMLM) such as PhotoActivated Localization Microscopy (PALM) [5], and Stochastic Optical Reconstruction (STORM) [48]. Some of these methods enhance image quality [8, 26, 43, 46], while others aim to upscale images by a factor [37, 47]. In contrast with standard SISR evaluations on photo-realistic images, where an LR image is typically a synthetic interpolated version of an HR image, SISR methods for microscopy images deal with real LR images. This translates into a difficult task since real LR images are produced through a different unknown process than a deterministic interpolation which is relatively easy to learn. Despite the success of these methods, the availability of datasets remains an issue in microscopy, where datasets are private and inaccessible. This data unavailability hinders research progress in microscopy imaging and prevents leveraging the potential of deep SISR methods.
@@
Our work addresses the lack of public datasets for SISR in fluorescence microscopy. To the best of our knowledge, no public dataset exists for this modality. SR-CACO-2 is a new confocal fluorescence dataset proposed for bio-imaging with pairs of HR and (real) LR images for the SISR task. It is based on human epithelial cell line Caco-2 [34] (ATCC HTB-37). The imaged tiles (Fig.3) capture epithelial cells isolated from colon tissue with colorectal adenocarcinoma. Three different proteins are marked and imaged, yielding different views of a cell: Survivin (CELL0), E-cadherin or Tubulin (CELL1), and Histone H2B (CELL2). The dataset is composed of 22 tiles with the HR tiles measuring ${\leavevmode\nobreak\ 9,300\times 9,300}$ pixels. They are captured by laser scanning over a regular raster. Three different LR scales are provided (Fig. 1): /2, /4, and /8. The SR-CACO-2 dataset is designed for convenient evaluation of machine learning models – patches are cropped to a size ${512\times 512}$ from HR tiles, and their corresponding patches from all LR tiles at different scales. These patches are only comprised of regions of interest (ROI) (i.e., cells), while irrelevant regions with black backgrounds are discarded. This allows the collection of ${9,937}$ patches that can be directly used for the training and testing of deep SISR models.
@@
A large and challenging dataset, SR-CACO-2, for confocal fluorescence SISR microscopy is proposed. It captures an epithelial cell line derived from colorectal adenocarcinoma grown as 3D spheroids. SR-CACO-2 contains 22 tiles in HR format and three corresponding real LR versions /2, /4, and /8 covering three separate protein markers: Survivin, E-cadherin or Tubulin, and Histone H2B.
@@
(3) An extensive benchmarking study of 15 representative SISR methods is provided to assess the quality of super-resolved images. These results show that state-of-the-art SISR methods yield smooth images and fail to correctly produce accurate HR texture across all scales. The full SR-CACO-2 dataset (tiles and patches), code, and pretrained weights of methods are made public.
(4) Images produced by all SISR methods are analyzed to assess their efficiency in downstream biology task for cell object (nucleus) detection and segmentation. Although they provide poor visual quality, several methods achieve promising results on these tasks compared to LR and HR.
@@ 2 Related Work
1. Photo-realistic SISR: A large progress has been observed in SISR for photo-realistic application [70]. The proposed methods mostly deal with supervised SISR task. Commonly, the degradation process is known beforehand and often interpolation is used to synthesize LR images from HR images. In the following, we divide super-resolution (SR) existing works into four main families based on their framework [70] which differs in the way the SR image is upscaled based on the input LR image which is the main key problem in SISR task (Fig.2). In particular, each family differs in what type of upscaling is used and its location in the model.
@@
Pre-upsampling super-resolution. Initial works circumvented the difficulty of mapping LR to HR images by performing an initial upscaling step using standard algorithms such as interpolation which is then refined using deep learning models. SRCNN [14, 15], one of the pioneering works in SISR, firstly adopted such approach where bicubic interpolation is initially used to upscale the input image, followed by a convolutional neural network (CNN). This CNN is trained end-to-end where it takes the upscaled input and aims to refine it, in a supervised way, to yield a SR image as close as possible to the HR image. A major advantage of these methods is that they only need to focus on the refinement since a major part of the SISR has already been done by interpolation. Other works followed similar approach with differences in training loss or architecture design [29, 30, 58, 59]. For instance, VDSR [30] employs residual learning as loss where the model aims to learn the residuals between the bicubic and the HR images, in addition to using a deeper model than SRCNN. Other methods often rely on deep models, in particular using recursion such as DRCN [29], DRNN [58], and MemNet [59]. Despite the success of these methods, they are very computationally expensive and they require a lot of memory since these very deep models perform all their operations in HR image space. Subsequent works aim to find different approaches to bypass this computational cost while improving the quality of the produced SR image.
@@
Iterative up-and-down sampling super-resolution. This family leverages back-projection technique [28] suggested in  [61] as a way to improve SR methods. In particular, it is used in an iterative way to reinforce the mutual dependency LR-HR in image pairs. For instance, DBPN [20] uses consecutive up and down layers. A residual error is then added to current image prediction which is then feed to the next layer. The final layer concatenates all the previous predictions, then, followed by a CNN layer which yields the final prediction. A parallel work, SRFBN [39], uses a similar strategy with more dense skip connections. Such approach has been also used in video super-resolution [21] but it has not been adopted largely by the community.
@@
Progressive upsampling super-resolution. This family of methods tackles two common issues in SISR. Exiting works upscale the LR image at once which is difficult. A second issues is that each scale requires training one model. This category aims to build a single model for all scales, while scaling up the LR image gradually. This creates less cumbersome models for deployment, and also allows multi-step upscaling which facilitates more SISR task. For instance, LapSRN [33] and MS-LapSRN [32] adopted Laplacian pyramid SR network. It is based on a cascade of CNN that learns residual sub-bands of high resolution images, which is added to an interpolated version of the image. This is done repeatedly where the scale increases until reaching the final requested scale where the interpolated images has accumulated all the residuals. ProSR [67] follows a similar approach however it uses the same input for interpolation instead of intermediate resulting image. The methods under this family greatly reduce the learning difficulty by decomposing the task into multiple steps that support each others. However, training models with multi-stage scales remains difficult and vulnerable to instability.
@@
Post-upsampling super-resolution. These methods perform most of the computations in LR dimensions while placing learnable upsampling layers toward the end for minimal computations at HR dimensions. The aim is to reduce the training computational cost which is a well known bottleneck in SISR [16, 54]. SISR field is dominated by this framework. Several works have been proposed. FSRCNN [16] improved SRCNN [14] by changing its architecture from a pre-sampling method to post-sampling one. This is done by introducing a deconvolution layer at the output layer while learning an end-to-end mapping from LR to HR space. To furthermore reduce computation, they shrink feature maps towards the input then expand them back toward the last layers while using small filter size but more mapping layers. Different methods focused on improving residual learning for deep models as it has been shown to be beneficial for SISR task, including EDSR [41], RCAN [77], RDN [79], CARN [2], and MSRN [36]. In SRDenseNet [62], authors leverage more skip connections in a very deep network. Adversarial generative models have also been introduced in SRGAN [35] and ESRGAN [66]. Non-local aggregation method has been extensively used in image restoration. Under the assumption that similar patches frequently recur in a natural scene image, several SISR methods have been proposed to leverage this idea. IGNN [80] extends the Non-local self-similarity idea across scales using graphs. Attention-based models have been also developed for SISR task to improve features such as RNAN [78], SAN [12], HAN [45], SwinIR [40], NLSN [42], ENLCN [73]. In GRL method [38], authors propose to explicitly model image hierarchies in the global, regional, and local range into a new self-attention module. In ACT method [74], authors improve self-attention by including local features from CNNs and long-range multi-scale dependencies captured by transformers. Omni-SR method leverages both spatial and channel-wise self-attention [65]. Recently, iterative-based approaches such as diffusion methods have attracted the community attention [49, 75]. However, they remain computationally expensive.
@@
2. Microscopy SISR: Compared to photo-realistic field, microscopy domain has seen very limited work. Each existing work is specialized in one single imaging technique. Some of these methods enhance image quality [8, 26, 43, 46], while others aim to upscale images by a factor [37, 47]. In [8], authors propose task-assisted generative adversarial network (TA-GAN) which incorporates an auxiliary task such as segmentation which has shown to be better than unassisted methods. It has been used to convert confocal to STED images. Authors in [26] perform a similar task using GANs. Deep models haven used in [43] for STORM modality while PALM modality has been tackled in [46]. Additionally, recurrent neural networks (RNN) have been used to upscale images for STORM modality [37]. Authors in [47] propose deep Fourier channel attention network for SIM modality. All these works employ private real or synthetic datasets. Only the work of [47] provides a public dataset for SIM modality. Such lack of public data can contribute in slowing down research in microscopy SISR field. Our work is an effort to provide a public dataset for fluorescence microscopy SISR task for model’s training and evaluation.
@@
Figure 3: Methodology for capturing of the SR-CACO-2 dataset.
@@ 3 The SR-CACO-2 Dataset
SR-CACO-2 is a dataset suitable for designing and evaluating machine/deep learning methods that can perform SISR in fluorescence microscopy imaging. In particular, imaged tiles of size ${\sim 9k\times 9k}$ capture the human epithelial cell line Caco-2. These cells are a well-established model for studying mitotic spindle orientation and epithelial cell polarity. As they are cell lines, they can be cultured over long periods without the requirement for repeated re-isolation from tissue. They are also easily modified with lentivectors to overexpress or knock-down proteins of interest, or to introduce fluorescently-tagged proteins for use in live imaging.
@@
SR-CACO-2 is comprised of more than 9k real pairs of LR and HR patches. It is a representative dataset in this field since it is captured by a standard process used in fluorescence microscopy. In addition, three different upscaling scenarios are available, X2, X4, and X8 allowing a better study of the limit of SISR methods. The process of capturing our SR-CACO-2 dataset is described in Sec.3.1, while the experimental protocol described in Sec.3.2 allows preparing the dataset for SISR design and evaluation of SISR methods.
Figure 4: Pre-processing tiles to patches. First, the HR tile of cell CELL2 (the brightest) is binarized to allow localizing cells only (i.e. ROI). A sliding window of size ${512\times 512}$ is used to scan the entire tile with specific overlap. Only windows with enough cell mass are kept while the rest are discarded. The same coordinates of the preserved patches are used for CELL1 and CELL2 of HR tiles. Coordinates of LR patches X2, X4, and X8 are computed by scaling down the HR patch coordinates accordingly to the corresponding scale.
@@
Table 1: SR-CACO-2 dataset splits and relevant statistics.
3.1 Dataset Capture:

### v2

@@ Abstract
Confocal fluorescence microscopy is one of the most accessible and widely used imaging techniques for the study of biological processes at the cellular and subcellular levels. Scanning confocal microscopy allows the capture of high-quality images from thick three-dimensional (3D) samples, yet suffers from well-known limitations such as photobleaching and phototoxicity of specimens caused by intense light exposure, which limits its use in some applications, especially for living cells. Cellular damage can be alleviated by changing imaging parameters to reduce light exposure, often at the expense of image quality. Machine/deep learning methods for single-image super-resolution (SISR) can be applied to restore image quality by upscaling lower-resolution (LR) images to produce high-resolution images (HR). These SISR methods have been successfully applied to photo-realistic images due partly to the abundance of publicly available datasets. In contrast, the lack of publicly available data partly limits their application and success in scanning confocal microscopy. In this paper, we introduce a large scanning confocal microscopy dataset named SR-CACO-2 that is comprised of low- and high-resolution image pairs marked for three different fluorescent markers. It allows to evaluate the performance of SISR methods on three different upscaling levels (X2, X4, X8). SR-CACO-2 contains the human epithelial cell line Caco-2 (ATCC HTB-37), and it is composed of 2,200 unique images, captured with four resolutions and three markers, that have been translated in the form of 9,937 patches for experiments with SISR methods. Given the new SR-CACO-2 dataset, we also provide benchmarking results for 16 state-of-the-art methods that are representative of the main SISR families. Results show that these methods have limited success in producing high-resolution textures, indicating that SR-CACO-2 represents a challenging problem. The dataset is released under a Creative Commons license (CC BY-NC-SA 4.0), and it can be accessed freely. Our dataset, code and pretrained weights for SISR methods are publicly available: https://github.com/sbelharbi/sr-caco-2.
@@ 1 Introduction
Deep learning models have recently provided significant improvements in diverse image analysis tasks [26] and biosciences [43, 69]. In particular, they have allowed for considerable advances in single-image super-resolution (SISR) [92] where high-resolution (HR) images are restored from low-resolution (LR) photo-realistic images [52, 85, 95, 99]. However, the success of these models relies heavily on the availability of large-scale public datasets to train a deep SISR model. Few works in microscopy imaging aim to leverage SISR [8, 15, 35, 51, 60, 63, 65]. This extends to different microscopy techniques [79] including standard fluorescence microscopy such as confocal [76], Structured Illumination Microscopy (SIM) [27], STimulated Emission Depletion (STED) microscopy [31], Single-Molecule Localization Microscopy (SMLM) such as PhotoActivated Localization Microscopy (PALM) [5], and Stochastic Optical Reconstruction (STORM) [66]. Some of these works enhance image quality [8, 35, 60, 63], while others aim to upscale images by a factor [51, 65]. In contrast with standard SISR evaluations on photo-realistic images, where an LR image is typically a synthetic downsampled version of an HR image, SISR models for microscopy images deal with real LR images. This translates into a difficult task since real LR images are produced through a different unknown process than a deterministic interpolation which is relatively easy to learn. Despite the success of these SISR models, the availability of datasets remains an issue in microscopy, where datasets are private and inaccessible. This data unavailability hinders progress in microscopy imaging research and prevents leveraging the potential of deep SISR models.
@@
Our work addresses the lack of public datasets for SISR in fluorescence microscopy. To the best of our knowledge, no public dataset exists for this modality. SR-CACO-2 is a new confocal fluorescence dataset proposed for bio-imaging with pairs of HR and (real) LR images for SISR. It is based on human epithelial cell line Caco-2 [46] (ATCC HTB-37). The imaged tiles (see Fig. 3) capture epithelial cells isolated from colon tissue with colorectal adenocarcinoma. Three different proteins are marked and imaged, yielding different views of a cell: Survivin (CELL0), E-cadherin or Tubulin (CELL1), and Histone H2B (CELL2). The dataset is composed of 2,200 unique images forming 22 tiles with the HR tiles measuring ${\leavevmode\nobreak\ 9,300\times 9,300}$ pixels. They are captured by laser scanning over a regular raster. Three different LR scales are provided (Fig. 1): /2, /4, and /8 in addition to the HR scale. The SR-CACO-2 dataset is designed for convenient evaluation of machine learning models – patches are cropped to a size ${512\times 512}$ from HR tiles, and their corresponding patches from all LR tiles at different scales. These patches are only comprised of regions of interest (ROI) (i.e., cells), while irrelevant regions with black backgrounds are discarded. This allows for the collection of ${9,937}$ patches, per scale and cell type, that can be directly employed for the training and testing of deep SISR models.
@@
A large, diverse, and challenging dataset, SR-CACO-2, for confocal fluorescence SISR microscopy is proposed. It captures an epithelial cell line derived from colorectal adenocarcinoma grown as 3D spheroids. SR-CACO-2 contains 2,200 unique images in HR format and three corresponding real LR versions /2, /4, and /8 covering three separate protein markers: Survivin, E-cadherin or Tubulin, and Histone H2B.
@@
(3) An extensive benchmarking study on 16 representative SISR methods is provided to assess the quality of super-resolved images. These results show that state-of-the-art SISR methods yield smooth images and fail to correctly produce accurate HR textures across all scales. The full SR-CACO-2 dataset (tiles and patches), code, and pretrained weights of methods are made public. The dataset is freely available under a Creative Commons license (CC BY-NC-SA 4.0).
(4) Images produced by all SISR methods are analyzed to assess their efficiency in downstream biology tasks for cell object (nucleus) detection and segmentation. Although they provide poor visual quality, several methods achieve promising results on these tasks compared to LR and HR.
@@ 2 Related Work
1. Photo-realistic SISR: A large progress has been observed in SISR for photo-realistic application [92, 96]. Earlier to deep learning methods, SISR methods pursued different approaches [68]. For instance, image statistical methods leverage different priors about the image such as minimal local variation by exploiting total variation [23], maximum entropy [64], and other priors [22]. Another dominant and successful technique is example-based approach [92, 96] where an image is decomposed into patches, and the super-resolved image is obtained by upscaling each patch. These methods typically rely on dictionaries and sparse coding [82, 102].
The introduction of deep SISR [19, 20], changes they way to tackle this task. The proposed methods mostly deal with supervised SISR task. Commonly, the degradation process is known beforehand and often interpolation is used to synthesize LR images from HR images. In the following, we divide deep super-resolution (SR) existing works into four main families based on their framework [92] which differs in the way the SR image is upscaled based on the input LR image which is the main key problem in SISR task (Fig.2). In particular, each family differs in what type of upscaling is used and its location in the model.
@@
Pre-upsampling super-resolution. Initial works circumvented the difficulty of mapping LR to HR images by performing an initial upscaling step using standard algorithms such as interpolation which is then refined using deep learning models. SRCNN [19, 20], one of the pioneering works in SISR, first adopted such an approach where bicubic interpolation is initially used to upscale the input image, followed by a convolutional neural network (CNN). This CNN is trained end-to-end where it takes the upscaled input and aims to refine it, in a supervised way, to yield a SR image as close as possible to the HR image. A major advantage of these methods is that they only need to focus on the refinement since a major part of the SISR has already been done by interpolation. Other works followed a similar approach with differences in training loss or architecture design [41, 42, 77, 78]. For instance, VDSR [42] employs residual learning as loss where the model aims to learn the residuals between the bicubic and the HR images, in addition to using a deeper model than SRCNN. Other methods often rely on deep models, in particular using recursions such as DRCN [41], DRNN [77], and MemNet [78]. Despite the success of these methods, they are very computationally expensive and they require a lot of memory since these very deep models perform all their operations in HR image space. Subsequent works aim to find different approaches to bypass this computational cost while improving the quality of the produced SR image.
@@
Iterative up-and-down sampling super-resolution. This family leverages the back-projection technique [37] suggested in  [80] as a way to improve SR methods. In particular, it is used in an iterative way to reinforce the mutual dependency LR-HR in image pairs. For instance, DBPN [28] uses consecutive up and down layers. A residual error is then added to the current image prediction which is then fed to the next layer. The final layer concatenates all the previous predictions, then, followed by a CNN layer which yields the final prediction. A parallel work, SRFBN [54], uses a similar strategy with more dense skip connections. Such an approach has been also used in video super-resolution [29] but it has not been adopted largely by the community.
@@
Progressive upsampling super-resolution. This family of methods tackles two common issues in SISR. Exiting works upscale the LR image at once which is difficult. A second issue is that each scale requires training one model. This category aims to build a single model for all scales while scaling up the LR image gradually. This creates less cumbersome models for deployment and also allows multi-step upscaling which facilitates more SISR tasks. For instance, LapSRN [45] and MS-LapSRN [44] adopted the Laplacian pyramid SR network. It is based on a cascade of CNN that learns residual sub-bands of high-resolution images, which is added to an interpolated version of the image. This is done repeatedly where the scale increases until reaching the final requested scale where the interpolated images has accumulated all the residuals. ProSR [89] follows a similar approach however it uses the same input for interpolation instead of the intermediate resulting image. The methods under this family greatly reduce the learning difficulty by decomposing the task into multiple steps that support each others. However, training models with multi-stage scales remains difficult and vulnerable to instability.
@@
Post-upsampling super-resolution. These methods perform most of the computations in LR dimensions while placing learnable upsampling layers toward the end for minimal computations at HR dimensions. The aim is to reduce the training computational cost which is a well-known bottleneck in SISR [21, 73]. The SISR field is dominated by this framework. Several works have been proposed. FSRCNN [21] improved SRCNN [19] by changing its architecture from a pre-sampling method to a post-sampling one. This is done by introducing a deconvolution layer at the output layer while learning an end-to-end mapping from LR to HR space. To furthermore reduce computation, they shrink feature maps towards the input and then expand them back toward the last layers while using a small filter size but more mapping layers. Different methods focused on improving residual learning for deep models as it is beneficial for SISR tasks, including EDSR [56], RCAN [103], RDN [105], CARN [2], and MSRN [50]. In SRDenseNet [81], authors leverage more skip connections in a very deep network. Adversarial generative models have also been introduced in SRGAN [47] and ESRGAN [88]. Non-local aggregation method has been extensively used in image restoration. Under the assumption that similar patches frequently recur in a natural scene image, several SISR methods have been proposed to leverage this idea. IGNN [106] extends the Non-local self-similarity idea across scales using graphs. Attention-based models have been also developed for SISR task to improve features such as RNAN [104], SAN [17], HAN [62], SwinIR [55], NLSN [59], ENLCN [95]. In GRL method [52], authors propose to explicitly model image hierarchies in the global, regional, and local range into a new self-attention module. In the ACT method [99], authors improve self-attention by including local features from CNNs and long-range multi-scale dependencies captured by transformers. Omni-SR method leverages both spatial and channel-wise self-attention [85]. Recently, iterative-based approaches such as diffusion methods have attracted the community’s attention [67, 100]. However, they remain computationally expensive.
A recent line of research tackles the task of arbitrary-scale SR (ASSR) [16, 24, 30, 39, 53, 57, 98] where the aim is to design techniques that can upscale an image to any arbitrary scale, mainly by relying on neural fields [12, 14, 48] to represent continuous signals that can be sampled at arbitrary rates. For instance, in [16], authors employ local implicit neural representation to represent images in a continuous domain. In particular, they propose a Local Implicit Image Function (LIIF) method. ASSR is achieved by replacing standard fixed-upsampling techniques with a multilayer perceptron (MLP) used to query pixel value at any coordinate and any scale.
@@
2. Microscopy SISR: Compared to the photo-realistic field, the microscopy domain has seen very limited work. Each existing work is specialized in one single imaging technique. Some of these methods enhance image quality [8, 35, 60, 63], while others aim to upscale images by a factor [51, 65]. In [8], authors propose a task-assisted generative adversarial network (TA-GAN) which incorporates an auxiliary task such as segmentation which has shown to be better than unassisted methods. It has been used to convert confocal to STED images. Authors in [35] perform a similar task using GANs. Deep models haven used in [60] for STORM modality while PALM modality has been tackled in [63]. Additionally, recurrent neural networks (RNN) have been used to upscale images for STORM modality [51]. Authors in [65] propose a deep Fourier channel attention network for SIM modality. In [58], authors consider a GAN-based approach to SISR of cytopathological images. All these works employ private real or synthetic datasets. Only the work of [65] provides a public dataset for SIM modality. Such lack of public data can contribute in slowing down research in the microscopy SISR field. Our work is an effort to provide a public dataset for fluorescence microscopy SISR tasks for model training and evaluation.
It is worth mentioning that there is an emerging line of research in SISR for natural scene images that deal with real-world SISR [11, 13, 38, 49, 97]. Such methods avoid using simulated low-resolution images obtained by interpolation. Instead, real low-resolution images are used to deal with real-world scenarios.
@@
Figure 3: Methodology for capturing the SR-CACO-2 dataset follows these steps:
@@ 3 The SR-CACO-2 Dataset
SR-CACO-2 is a dataset suitable for designing and evaluating machine/deep learning methods that can perform SISR in fluorescence microscopy imaging. In particular, image tiles of size ${\sim 9k\times 9k}$ are comprised of ${10\times 10}$ unique images, and capture the human epithelial cell line Caco-2. These cells are a well-established model for studying mitotic spindle orientation and epithelial cell polarity. As they are cell lines, they can be cultured over long periods without the requirement for repeated re-isolation from tissue. They are also easily modified with lentivectors to overexpress or knock-down proteins of interest, or to introduce fluorescently-tagged proteins for use in live imaging.
@@
SR-CACO-2 is comprised of more than 9k real pairs of LR and HR patches, per scale, and cell type. It is a representative dataset in this field since it is captured using a standard process in fluorescence microscopy. In addition, three different upscaling scenarios are provided, X2, X4, and X8, in addition to HR, allowing a better study of the limit of SISR methods. The process of capturing our SR-CACO-2 dataset is described in Sec.3.1, while the experimental protocol described in Sec.3.8 allows preparing the dataset for the design and evaluation of SISR models.
Figure 4: Pre-processing of tiles to patches. First, the HR tile of cell CELL2 (the brightest) is binarized to allow localizing cells only (i.e. ROI). A sliding window of size ${512\times 512}$ is used to scan the entire tile with specific overlap. Only windows with enough cell mass are preserved while the rest are discarded. The same coordinates of the preserved patches are used for CELL1 and CELL2 of HR tiles. Coordinates of LR patches X2, X4, and X8 are computed by scaling down the HR patch coordinates according to the corresponding scale.
@@
Table 1: Split of SR-CACO-2 into the train, validation and test subsets, along with relevant statistics. Numbers are defined per scale (X2, X4, X8, and HR), and per cell type (CELL0, CELL1, and CELL2).
3.1 Dataset Capture